OpenStudy (anonymous):

Create your own two-step equation with All Real Numbers as the solution. Show the steps to solving the equation to prove the solution All Real Numbers.

OpenStudy (rulnick):

3x+1 = 1 - (-3x)

OpenStudy (rulnick):

sub 1 from both sides 3x = -(-3x)

OpenStudy (rulnick):

or 3x = 3x divide both sides by 3

OpenStudy (rulnick):

x=x all real numbers satisfy this

OpenStudy (anonymous):

could you write it all in one reply it's confusing me.

OpenStudy (rulnick):

3x+1 = 1 + 3x -1 -1 ------ ----------- 3x = 3x /3 /3 ------- -------- x = x

OpenStudy (anonymous):

Still so confused >.<

OpenStudy (anonymous):

if you are left with a true statement when solving for a variable that is an identity such as x=x, or evaluating this expression 0=0, then the solution set is all real numbers

OpenStudy (anonymous):

I don't understand the example though, Like how would I write it out?

OpenStudy (anonymous):

he already wrote it out in his example, there are many different examples you can use